OpenStudy (anonymous):

Best answer gets a medal (: Simplify. -14x^3 / x^3-5x^4 A. -14/5x-1 B. -14x/1-5x C. 1-5x/-14x D. -14/1-5x

OpenStudy (anonymous):

You can first take a factor of x^3 out from the top and the bottom and they will cancel.\[\frac{ -14x^3 }{ x^3 -5x^4 }\] \[\frac{ (x^3)(-14) }{ (x^3)(1-5x) }\]and the x^3 cancels to give you this. \[\frac{ -14 }{ 1-5x }\]

OpenStudy (anonymous):

Thank You So Much ! (: